I need help solving this problem 3x-7=x-9
Found 2 solutions by checkley75, heidi:
Answer by checkley75(3666)

3X-7=X-9
3X-X=7-9
2X=-2
X=-2/2
X=-1 ANSWER.
PROOF
3*-1-7=-1-9
-3-7=-10
-10=-10

Answer by heidi(17)

3x-7 = x-9
+7 +7
3x=-2
+2 +2
5x